A potter's wheel having a radius 0.49 m and a moment of inertia of 12.4 kg · m2 is rotating freely at 46 rev/min. The potter can stop the wheel in 5.5 s by pressing a wet rag against the rim and exerting a radially inward force of 71 N. Find the effective coefficient of kinetic friction between the wheel and the wet rag.
Answer #1

ω = ω0 + α t.
ω0 = 2π *46/60, ω = 0 and t = 5.5 s.
0 = 2π *46/60 - 5.5α
solve for α
α = 0.875 rad/s^2

Torque = Moment of Inertia * Angular acceleration
Torque = 12.4 kg m^2 * 0.875 rad/s^2
Torque = 10.85 N-m

τ = r * f
r * μ * F = 10.85
0.49 * μ * 71 = 10.85
μ = 0.312

Coefficient of Kinetic friction, μ =  0.312
a = r α
μ = │a│/g